OpenStudy (anonymous):

s=5r^2t solve for T please

OpenStudy (bahrom7893):

t=s/(5r^2)

OpenStudy (bahrom7893):

just divide both sides by 5r^2

OpenStudy (anonymous):

log s=2t log 5r log s/(2*log 5r)=T T=log(s-5r)/2 @ bahrom shoudnt it be this??

OpenStudy (anonymous):

check it @bahrom7893

OpenStudy (bahrom7893):

That's right sarkar